# Splitting differential equations using Galois theory

### v1

@@ 2.2. Definable Jordan decomposition
In this subsection, we take inspiration from the definable Chevalley and Rosenlicht decompositions of [6, Subsection 2.3] and produce a definable Jordan decomposition for groups isomorphic to an algebraic group. For this subsection only, we work with some arbitrary theory $\mathrm{Th}$ eliminating imaginaries (we do not call it $T$ to avoid confusion with tori, which we will denote $T$), some sufficiently saturated $\mathcal{U}\models\mathrm{Th}$, and some small algebraically closed sets of parameters $A\subset B\subset\mathcal{U}$. Let $\mathcal{C}$ be an $A$-definable purely stably embedded algebraically closed field.
In this subsection, we will extend the Jordan decomposition for linear solvable algebraic groups to groups definably isomorphic to the $\mathcal{C}$-points of a linear solvable algebraic group. In particular, we prove the existence of a definable unipotent radical, as well as some basic properties. Note that our methods would in fact allow us to define the definable unipotent radical of any group definably isomorphic to the $\mathcal{C}$-points of a linear algebraic group.
Recall that a torus is an algebraic group that is isomorphic to $G^{n}_{m}$, for some $n$. If a torus is defined over some algebraically closed field $F$, then this isomorphism is defined over $F$ (see [4, 34.3]). We call a group definably isomorphic to the $\mathcal{C}$-points of a torus a definable torus, and use a similar terminology for definable unipotent, nilpotent and solvable linear algebraic groups.
@@
We have the following, which is a consequence of the Lie-Kolchin Theorem (see [4, Chapter 19] for a proof):
@@
Let $G$ be a solvable $B$-definable group, and $f:G\rightarrow\widehat{G}(\mathcal{C})$ be a definable isomorphism, where $\widehat{G}$ is a connected linear algebraic group. Then there is a $B$-definable normal subgroup $G_{u}\trianglelefteq G$ which is the unique maximal subgroup of $G$ definably isomorphic to the $\mathcal{C}$-points of a unipotent group. We thus obtain an exact sequence:
@@
and $T:=G/G_{u}$ is definably isomorphic to the $\mathcal{C}$ points of a torus.
The proof is similar to the proof of [6, Fact 2.8]:
@@
By Fact 2.14 we have a definable, definably split, short exact sequence $1\rightarrow\widehat{G}_{u}\rightarrow\widehat{G}\rightarrow\widehat{T}\rightarrow 1$. Let $G_{u}=f^{-1}(\widehat{G}_{u}(\mathcal{C})).$ This is a definable subgroup of $G$ which is definably isomorphic to $\widehat{G}_{u}(\mathcal{C}),$ the $\mathcal{C}$ points of the unipotent group $\widehat{G}_{u}$.
To show that $G_{u}$ is unique and maximal, let $H\leq G$ be another definable subgroup, and let $l$ be a definable isomorphism such that $l(H)=\widehat{H}(\mathcal{C})$, the $\mathcal{C}$-points of an unipotent linear algebraic group $\widehat{H}$. Since $\mathcal{C}$ is purely stably embedded, the map $f\circ l^{-1}$ is a morphism of algebraic groups, and thus $f(H)$ is an unipotent algebraic subgroup of $\widehat{G}(\mathcal{C})$. By Fact 2.14, it must be contained in $\widehat{G}_{u}(\mathcal{C})$, and therefore $H\leq G_{u}$.
@@
Finally, the linear algebraic group $\widehat{T}$ is a torus, and $T$ is definably isomorphic to $\widehat{T}(\mathcal{C})$.
We call this normal subgroup the definable unipotent radical of $G$. We show that, just as is the case for linear algebraic groups, it is preserved by definable maps:
@@
We obtain, by stable embeddedness of $\mathcal{C}$, a $\mathcal{C}\cap B$-definable morphism of linear algebraic groups $\widehat{\pi}=h\circ\pi\circ f^{-1}:\widehat{G}(\mathcal{C})\rightarrow\widehat{K}(\mathcal{C})$, and we must have, by [4, Section 15.3] that $\widehat{\pi}(\widehat{G}_{u})<\widehat{K}_{u}$. Applying $h^{-1}$, we obtain $\pi\circ f^{-1}(\widehat{G}_{u})=h^{-1}(\widehat{K}_{u})$. By construction of the definable unipotent radical, this gives us $\pi(G_{u})<K_{u}$. ∎
@@
Here $\widehat{G}$ is a nilpotent connected linear algebraic group, and thus by [4, Proposition 19.2] splits into a product $\widehat{G_{u}}\times\widehat{T}$ for some unipotent $\widehat{G_{u}}$ and torus $\widehat{T}$. These groups are the maximal unipotent subgroup and torus of $\widehat{G}$. Using this, we can prove, as was done in Lemma 2.15, that their preimages under $f$ are normal $B$-definable subgroups of $G$, and it then follows that $G$ is $B$-definably isomorphic to $G_{u}\times T$. ∎
@@
where $K$ is a definable torus, and $G/K$ is definably nilpotent. Then $G$ is nilpotent and $\widehat{G}$ is linear. If moreover the maximal tori of $G$ and $G/K$ are $B$-definably isomorphic to the $\mathcal{C}$-points of tori, then the short exact sequence is $B$-definably split and $G=K\times G/K$.
@@
As an extension of a solvable group by an abelian group, the group $G$ is solvable, and it is linear as an extension of a linear group by a linear group. By [4, Proposition 19.4], as $K$ is a definable torus, this implies that $C_{G}(K)=N_{G}(K)$, and as $K$ is normal, is must be central. Therefore $G$ is nilpotent because it is a central extension of a nilpotent group.
By Lemma 2.17, we know that $G$ is $B$-definably isomorphic to the product of its maximal unipotent subgroup and torus, and the same is true for $G/K$. Consider the following diagram, where all vertical and horizontal short sequences are exact:
{tikzcd}
@@
By assumption, the maximal tori $T$ and $S$ are $B$-definably isomorphic to the $\mathcal{C}$-points of tori $\widehat{T}$ and $\widehat{S}$, which are by stable embeddedness defined over $\mathcal{C}\cap B$. Using the machinery of characters of $d$-groups (see [4, Section 16.2]), we can prove that the induced map $\widehat{T}\rightarrow\widehat{S}$ is split. Moreover, because $B\cap\mathcal{C}$ is algebraically closed, by [25, Proposition 3.2.12], it is $B\cap\mathcal{C}$-definably split. Therefore the induced short exact sequence from $T$ to $S$ is $B$-definably split, implying that $T$ is $B$-definably isomorphic to $K\times S$.
Thus we have obtained that $G$ and $G/K$ are $B$-definably isomorphic to $G_{u}\times S\times K$ are $G_{u}\times S$, respectively. We can conclude immediately from this.
@@ 3.1. Definable fibrations and uniform internality
In this section, we go back to considering a stable theory $T$ eliminating imaginaries, and further restrict to $T$ being $\omega$-stable, as to have definable binding groups. We fix an algebraically closed set of parameters $A$, and some $A$-definable set $\mathcal{C}$.
If $p\in S(A)$, by an $A$-definable map $f:p\rightarrow f(p)$, what we mean is an $A$-definable map $f$ with domain containing $p$, which restricts into a relatively definable map on $p(\mathcal{U})$. The image $f(p(\mathcal{U}))$ is then the set of realisations of a type over $A$, which we denote $f(p)$.
The question we will try to answer concerns when a type is interdefinable, or interalgebraic, with a product of types. Of course, given a type $p$, it is always interdefinable with $p\otimes q$, where $q$ is a type with only one realization. To make the question interesting, we ask the type to split along a map, and more precisely a definable fibration:
@@
Let $p\in S(A)$ be stationary. An $A$-definable map $f:p\rightarrow f(p)$ is a fibration if for any $a\models p$, the type $\operatorname{tp}(a/f(a)A)$ is stationary.
Note that the set of realizations of $\operatorname{tp}(a/f(a)A)$ is exactly the fiber $f^{-1}(\{a\})$. We introduce a notation for these types:
@@
If $p\in S(A)$ and $f:p\rightarrow f(p)$ is an $A$-definable map, then for any $f(a)\models f(a)$, we denote $p_{f(a)}:=\operatorname{tp}(a/f(a)A)$.
More specifically, we will mostly be interested in fibrations such that each fiber $p_{f(a)}$ is almost internal to some fixed definable set $\mathcal{C}$. These are called relatively internal fibrations in [5].
Now suppose that the type $f(p)$ is almost $\mathcal{C}$-internal. When is $p$ almost $\mathcal{C}$-internal as well? The answer is given by uniform internality, introduced in [5]:
@@
Let $p\in S(A)$ be stationary, and $f:p\rightarrow f(p)$ a fibration. The fibration $f:p\rightarrow f(p)$ is said to be uniformly $\mathcal{C}$-internal (resp. uniformly almost $\mathcal{C}$-internal) if the fibers $p_{f(a)}$ are $\mathcal{C}$-internal (resp. almost $\mathcal{C}$-internal) and there is a tuple $e$ such that for some (any) $a\models p$, independent from $e$ over $A$, we have that $a\in\operatorname{dcl}(e,f(a),\mathcal{C})$ (resp. $\operatorname{acl}$).
The following is easy to prove (see [6, Proposition 3.16]):
@@
Let $p\in S(A)$ be stationary, and $f:p\rightarrow f(p)$ a fibration with almost $\mathcal{C}$-internal fibers. The type $p$ is almost $\mathcal{C}$-internal if and only if $f(p)$ is almost $\mathcal{C}$-internal and $f$ is uniformly almost $\mathcal{C}$-internal.
If $p$ is interalgebraic with $f(p)\otimes s$, for some almost $\mathcal{C}$-internal $s\in S(A)$, then it is easy to prove that $f$ is uniformly almost $\mathcal{C}$-internal. In the rest of this article, we will be preoccuppied with the converse: given an uniformly internal map, is it reasonable to expect it to come from a product? This was asked directly in [5] and indirectly in [10]. We start by giving a precise definition of splitting.
@@
Graphically, splitting means that there is an $A$-definable bijection between $p$ and $f(p)\otimes r$ such the the diagram:
{tikzcd}
commutes. When the fibration is clear from context we may write about (almost) split types. Up to interalgebraicity, almost split and split are the same:
Remark 3.6.
Let $f:p\rightarrow f(p)$ be an almost split $A$-definable fibration, and let $r$ be the type witnessing it. Then $p$ is interalgebraic with $f(p)\otimes r$. Moreover, the fibration $\pi:f(p)\otimes q\rightarrow f(p)$ has almost internal fibers, and is split. Thus any almost split type is interalgebraic with a split type.
Using this remark, we can also reduce, up to interalgebraicity, to a split fibration with internality replacing almost internality. Recall the following (see [10, Lemma 3.6] for a proof):
Fact 3.7.
If $p=\operatorname{tp}(a/A)$ is an almost $\mathcal{C}$-internal type, then there is some $b\in\operatorname{dcl}(aA)$ such that $a\in\operatorname{acl}(bA)$ and $\operatorname{tp}(b/A)$ is $\mathcal{C}$-internal. In particular, any almost $\mathcal{C}$-internal type is interalgebraic with an internal type.
Using Remark 3.6 and applying Fact 3.7 to $q$, we get:
@@
As mentioned previously, it is straightforward to show that splitting implies uniform internality (see for example, [5, Proposition 3.10]):
Proposition 3.9.
If $f:p\rightarrow f(p)$ is an almost split $A$-definable fibration with almost $\mathcal{C}$-internal fibers, then it is uniformly almost $\mathcal{C}$-internal.
It is natural to ask about the converse, as splitting is, a priori, a much stronger condition. This is one of the themes of [5], which focuses on the case where $f(p)$ is orthogonal to $\mathcal{C}$. In that case, splitting and uniform internality coincide:
Proposition 3.10 ([5], Proposition 3.14).
Let $f:p\rightarrow f(p)$ be a fibration with almost $\mathcal{C}$-internal fibers. If $f(p)$ is orthogonal to $\mathcal{C}$, then $f$ is almost split if and only if it is uniformly almost $\mathcal{C}$-internal.
Therefore if we are interested in whether or not a uniformly almost $\mathcal{C}$-internal $f:p\rightarrow f(p)$ splits, we may as well assume that $f(p)\not\perp\mathcal{C}$. In this article, we will be interested in the case where $f(p)$ is itself $\mathcal{C}$-internal.

### v3

@@ 2.2. Definable Jordan decomposition
In this subsection, we take inspiration from the definable Chevalley and Rosenlicht decompositions of [8, Subsection 2.3] and produce a definable Jordan decomposition for groups isomorphic to an algebraic group. For this subsection, we pick algebraically closed sets of parameters $A\subset B$, and we assume that $\mathcal{C}$ is an $A$-definable purely stably embedded algebraically closed field.
We will extend the Jordan decomposition for linear solvable algebraic groups to groups definably isomorphic to the $\mathcal{C}$-points of a linear solvable algebraic group. In particular, we prove the existence of a definable unipotent radical, as well as some basic properties.
Recall that a torus is an algebraic group that is isomorphic to $G^{n}_{m}$, for some $n$, and a diagonalizable group is a group isomorphic to a definable subgroup of $G^{n}_{m}$ for some $n$. If a torus is defined over some algebraically closed field $F$, then this isomorphism is defined over $F$ (see [6, 34.3]). A unipotent group is an algebraic group that is isomorphic to an algebraic subgroup of $U_{n}$, the group of $n\times n$ upper-triangular matrices with $1$’s along the diagonal.
We call a group definably isomorphic to the $\mathcal{C}$-points of a torus a definable torus, and use a similar terminology for definable unipotent, nilpotent, solvable linear algebraic groups and diagonalizable groups.
@@
We recall the multiplicative Jordan decomposition, which is a consequence of the Lie-Kolchin Theorem (see [6, Chapter 19] for a proof):
@@
Let $G$ be a solvable $B$-definable group, and $f:G\rightarrow\widehat{G}(\mathcal{C})$ be a definable isomorphism, where $\widehat{G}$ is a connected linear algebraic group. Then there is a $B$-definable normal subgroup $G_{u}\trianglelefteq G$ which is the unique maximal subgroup of $G$ definably isomorphic to the $\mathcal{C}$-points of a unipotent group. We thus obtain a definable exact sequence:
@@
which is definably (maybe over extra parameters) split, and $T:=G/G_{u}$ is definably isomorphic to the $\mathcal{C}$ points of a torus.
The proof is similar to the proof of [8, Fact 2.8]:
@@
By Fact 2.14 we have a definable, definably split, short exact sequence $1\rightarrow\widehat{G}_{u}\rightarrow\widehat{G}\rightarrow\widehat{T}\rightarrow 1$. Let $G_{u}=f^{-1}(\widehat{G}_{u}(\mathcal{C}))$, then the sequence $1\rightarrow G_{u}\rightarrow G\rightarrow G/G_{u}\rightarrow 1$ is definably isomorphic to the previous sequence of algebraic groups, and thus also definably split. The group $G_{u}$ is a definable subgroup of $G$ which is definably isomorphic to $\widehat{G}_{u}(\mathcal{C}),$ the $\mathcal{C}$-points of the unipotent group $\widehat{G}_{u}$.
To show that $G_{u}$ is unique and maximal, let $H\leq G$ be another definable subgroup, and let $l$ be a definable isomorphism such that $l(H)=\widehat{H}(\mathcal{C})$, the $\mathcal{C}$-points of an unipotent linear algebraic group $\widehat{H}$. Since $\mathcal{C}$ is purely stably embedded, the map $f\circ l^{-1}$ is a morphism of algebraic groups, and thus $f(H)$ is a unipotent algebraic subgroup of $\widehat{G}(\mathcal{C})$. By Fact 2.14, it must be contained in $\widehat{G}_{u}(\mathcal{C})$, and therefore $H\leq G_{u}$.
@@
Finally, the algebraic group $\widehat{T}$ is a torus, and $T$ is definably isomorphic to $\widehat{T}(\mathcal{C})$. ∎
We call this subgroup the definable unipotent radical of $G$. We show that, just as is the case for linear algebraic groups, it is preserved by definable maps:
@@
We obtain, by stable embeddedness of $\mathcal{C}$, a $\mathcal{C}\cap B$-definable morphism of linear algebraic groups $\widehat{\pi}=h\circ\pi\circ f^{-1}:\widehat{G}(\mathcal{C})\rightarrow\widehat{K}(\mathcal{C})$, and we must have, by [6, Section 15.3] that $\widehat{\pi}(\widehat{G}_{u})<\widehat{K}_{u}$. Applying $h^{-1}$, we obtain $\pi\circ f^{-1}(\widehat{G}_{u})=h^{-1}(\widehat{K}_{u})$. By construction of the definable unipotent radical, this gives us $\pi(G_{u})<K_{u}$. ∎
@@
Here $\widehat{G}$ is a nilpotent connected linear algebraic group, and thus by [6, Proposition 19.2] splits into a product $\widehat{G_{u}}\times\widehat{T}$ for some unipotent $\widehat{G_{u}}$ and torus $\widehat{T}$. These groups are the maximal unipotent subgroup and torus of $\widehat{G}$. Using this, we can prove, as was done in Lemma 2.15, that their preimages under $f$ are normal $B$-definable subgroups of $G$, and it then follows that $G$ is $B$-definably isomorphic to $G_{u}\times T$. ∎
@@
where $K$ is a definable diagonalizable group, and $G/K$ is definably nilpotent linear. Then $G$ is nilpotent and $\widehat{G}$ is linear. If moreover $K$ is a torus and the maximal tori of $G$ and $G/K$ are $B$-definably isomorphic to the $\mathcal{C}$-points of tori, then the short exact sequence is $B$-definably split and $G=K\times G/K$.
@@
As an extension of a solvable group by an abelian group, the group $G$ is solvable, and $\widehat{G}$ is linear as an extension of a linear group by a linear group. By [6, Proposition 19.4], as $K$ is a definable diagonalizable group, this implies that $C_{G}(K)=N_{G}(K)$, and as $K$ is normal, is must be central. Therefore $G$ is nilpotent because it is a central extension of a nilpotent group.
Now suppose that $K$ is a torus. By Lemma 2.17, we know that $G$ is $B$-definably isomorphic to the product of its maximal unipotent subgroup and torus, and the same is true for $G/K$. Consider the following diagram, where all vertical and horizontal short sequences are exact and the vertical sequences are given by Lemma 2.17:
[…]
@@
By assumption, the maximal tori $T$ and $S$ are $B$-definably isomorphic to the $\mathcal{C}$-points of tori $\widehat{T}$ and $\widehat{S}$, which are by stable embeddedness defined over $\mathcal{C}\cap B$. Using the machinery of characters of $d$-groups (see [6, Section 16.2]), we can prove that the induced map $\widehat{T}\rightarrow\widehat{S}$ is split. Moreover, because $B\cap\mathcal{C}$ is algebraically closed, by [29, Proposition 3.2.12], it is $B\cap\mathcal{C}$-definably split. Therefore the induced short exact sequence from $T$ to $S$ is $B$-definably split, implying that $T$ is $B$-definably isomorphic to $K\times S$.
Thus we have obtained that $G$ and $G/K$ are $B$-definably isomorphic to $G_{u}\times S\times K$ are $G_{u}\times S$, respectively. We can conclude immediately from this. ∎
We conclude this section with a lemma on decomposition of actions of definably nilpotent linear algebraic groups:
Lemma 2.19.
Let $G=G_{u}\times T$ be a definable nilpotent linear algebraic group, where $G_{u}$ is its unipotent radical and $T$ its maximal torus. Consider a definable group action of $G$ on some definable set $X$. Then for all $a\in X$, we have $G_{u}a\cap Ta=\{a\}$.
Proof.
Let $\pi_{1}$ and $\pi_{2}$ be the projections on $G_{u}$ and $T$. We first prove:
Claim.
Proof.
By Lemma 2.9 applied to the group $\pi_{1}(H)\times\pi_{2}(H)$, there is a definable isomorphism between $\pi_{1}(H)/N_{1}$ and $\pi_{2}(H)/N_{2}$, where $N_{1}=\ker(\pi_{2}|_{H})$ and $N_{2}=\ker(\pi_{1}|_{H})$ are normal definable subgroups of $H$, which we can identify with subgroups of $\pi_{1}(H)$ and $\pi_{2}(H)$. But the group $\pi_{1}(H)/N_{1}$ is definably unipotent and the group $\pi_{2}(H)/N_{2}$ is a quotient of a definable torus, so there cannot be any definable isomorphism between them, unless they are both trivial. Therefore $\pi_{1}(H)=N_{1}$ and $\pi_{2}(H)=N_{2}$, which is equivalent to $H=\pi_{1}(H)\times\pi_{2}(H)$. ∎
The action of $G$ on $X$ is definably isomorphic to its action on $G/H$, where $H<G$ is a definable subgroup. By the claim, we know that $H=\pi_{1}(H)\times\pi_{2}(H)$, and thus the action is definably isomorphic to the action of $G=G_{u}\times T$ on $G_{u}/\pi_{1}(H)\times T/\pi_{2}(H)$. In particular for any $a\in X$, we have $G_{u}a\cap Ta=\{a\}$. ∎
Remark 2.20.
We have not used that the ambient theory is $\omega$-stable in this subsection, and the reader can check that our results go through in any theory eliminating imaginaries, as long as $\mathcal{C}$ is a purely stably embedded algebraically closed field.
@@ 3.1. Definable fibrations and uniform internality
In this section, we are still working in an $\omega$-stable theory $T$ eliminating imaginaries. We fix some algebraically closed set of parameters $A$, and some $A$-definable set $\mathcal{C}$.
If $p\in S(A)$, by an $A$-definable map $f:p\rightarrow f(p)$, what we mean is an $A$-definable map $f$ with domain containing $p$, which restricts into a relatively definable map on $p(\mathcal{U})$. The image $f(p(\mathcal{U}))$ is then the set of realisations of a type over $A$, which we denote $f(p)$. We will mainly be concerned with definable fibrations:
@@
Let $p\in S(A)$. An $A$-definable map $f:p\rightarrow f(p)$ is a fibration if for any $a\models p$, the type $\operatorname{tp}(a/f(a)A)$ is stationary.
Note that the set of realizations of $\operatorname{tp}(a/f(a)A)$ is exactly the fiber $f^{-1}(\{f(a)\})\cap p(\mathcal{U})$. We introduce a notation for these types:
@@
If $p\in S(A)$ and $f:p\rightarrow f(p)$ is an $A$-definable map, then for any $f(a)\models f(p)$, we denote $p_{f(a)}:=\operatorname{tp}(a/f(a)A)$.
More specifically, we will mostly be interested in fibrations such that each fiber $p_{f(a)}$ is almost internal to some fixed definable set $\mathcal{C}$. These are called relatively internal fibrations in [7]. We are interested in $\mathcal{C}$-internality of $p$, under the assumption that $f(p)$ is $\mathcal{C}$-internal. An answer is given by uniform internality, introduced in [7]:
@@
Let $p\in S(A)$, and $f:p\rightarrow f(p)$ a fibration. The fibration $f:p\rightarrow f(p)$ is said to be uniformly $\mathcal{C}$-internal (resp. uniformly almost $\mathcal{C}$-internal) if there is a tuple $e$ such that for some (any) $a\models p$, independent from $e$ over $A$, we have that $a\in\operatorname{dcl}(e,f(a),\mathcal{C})$ (resp. $\operatorname{acl}$).
This implies that the fibers $p_{f(a)}$ are (almost) $\mathcal{C}$-internal, but is strictly stronger. The following is easy to prove (see [8, Proposition 3.16]):
@@
Let $p\in S(A)$, and $f:p\rightarrow f(p)$ a fibration with almost $\mathcal{C}$-internal fibers. The type $p$ is almost $\mathcal{C}$-internal if and only if $f(p)$ is almost $\mathcal{C}$-internal and $f$ is uniformly almost $\mathcal{C}$-internal.
In this article, we are interested in a strictly stronger condition implying internality, which we call splitting (and is called triviality in [7]):
@@
Graphically, splitting means that there is an $A$-definable bijection between $p$ and $f(p)\otimes r$ such that the diagram:
${p}$${f(p)\otimes r}$${f(p)}$$\scriptstyle{f}$
commutes. It is key to our methods to be able to replace almost internal types with internal ones. To do so, the following is useful (see [12, Lemma 3.6] for a proof):
Fact 3.6.
If $p=\operatorname{tp}(a/A)$ is an almost $\mathcal{C}$-internal type, then there is some $b\in\operatorname{dcl}(aA)$ such that $a\in\operatorname{acl}(bA)$ and $\operatorname{tp}(b/A)$ is $\mathcal{C}$-internal. In particular, any almost $\mathcal{C}$-internal type is interalgebraic with a $\mathcal{C}$-internal type.
Applying Fact 3.6, it is easy to prove: